Thanks for helping me out! Here is the question? In 1770, Thomas Gainsborough painted The blue Boy. The sum of the length and width of the painting is 118 inches. The difference of the length and width is 22 inches. Find the lengh and width.
Answer by AlgebraLady88(44)

The sum of the length and width is 118
L+W = 118
The difference of the length and width is 22
L-W = 22
Now , what we have here is a 2 by 2 equation.
We can solve it nicely by adding up the two.
L+W= 118
L-W= 22
Adding the two equations will give us 2L = 140
L= 70
Plug L=70 into any of the two equations. We will use the first one
L+ W= 118
70 + W= 118
W= 118-70
W= 48
So, length is 70 and width is 48.